Dal and Kim can assemble a swing set in 1 1/2 hr. Working alone, it takes Kim 4 hr longer than Dal to assemble the swing set. How long would it take Dal, working alone, to assemble the swing set? I do not have a clue how to set this up Please help!!!
Answer by mananth(16949)

Dal and Kim can assemble a swing set in 1 1/2 hr. Working alone, it takes Kim 4 hr longer than Dal to assemble the swing set. How long would it take Dal, working alone, to assemble the swing set?
..
let Dal assemble in x hours
so he does 1/x of the job in 1 hour
..
kim takes (x+4) hours
so he does 1/(x+4) of the job in 1 hour.
..
Together they assemble in 3/2 hours
..
1/x + 1/(x+4) = 3/2
(x+4+x)/x(x+4)= 3/2
2x+4 /x(x+4)= 3/2
2(2x+4)=3x(x+4)
4x+8 = 3x^2+12x
3x^2-4x+12x-8 =0
3x^2+8x-8=0
..
find the roots by quadratic equation
discriminant = 160
x1= (-8+sqrt(160))/6
0.77 hours
Dal takes 0.77 hours to complete the job alone
...
m.ananth@hotmail.ca